HUBBLE SPACE TELESCOPE DAILY REPORT #844 PERIOD COVERED: 0800 EST Daily Status Report as of 075/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/RD 4112 (Evolution of Alpha Lyman and C IV Absorption Systems) The Faint Object Spectrograph was used to make observations of Lyman Alpha and Carbon IV emissions in quasar PG1216+069. The observations were executed as scheduled, and no problems were reported. 1.2 Completed PC 4714 (WF/PC Calibration: PC Internal Delta Flats) The Planetary Camera was used to make six internal flat observations which will be used for the restoration of Earth flats, which were altered due to changes in quantum efficiency which occurred after WFPC decontamination operations. All six observations were reported to look normal. 1.3 Completed FOC/96 4804 (UV Imaging of Nearby Galaxies with FOC/96) The Faint Object Camera (f/96) was used to to take ultraviolet snapshots of nearby galaxy UGC-77212. Three observations were monitored. On one of the observations there was data receipt problems. Too much of the data was missing to evaluate this observation. It was received okay by the Data Capture Facility (DCF). 1.4 Completed GHRS/2 3496 (Ultraviolet Atlas of Sirius) The High Resolution Spectrograph was used to attempt an ultraviolet spectral observation of the star Sirius (aka, HD48915). The baseline guide star acquisition for this series failed. The science observations were not executed. 1.5 Completed FOS 4698 (FOS Dark Monitoring) The Faint Object Spectrograph was used to make a series of dark calibration images on both the red and blue side to determine the instruments background. The dark exposures were monitored by the OSS, and no problems were seen. One observations showed a little bump in the middle of the data. The SHP (Standard Header Packet) and UDL (Unique Data Log) files for one observations were lost due to a science tape recorder track change. 1.6 Completed PC 4249 (Black Holes, Stellar Dynamics and Populations) The Planetary Camera was used to obtain four images of the nuclei of the large elliptical galaxy NGC-4623 which is located in the Virgo Cluster. All four observations were monitored and displayed in the OSS. Three of the observations showed some recentering during their execution. 1.7 Completed PC 4712 (WF/PC Calibration Measles Monitor) The Planetary Camera was used to make three Earth flat calibration observations using filter F517N, both with and without the filter F122M. All three observations were received and displayed in the OSS. The first and third observations were reported to be saturated, and the second observations was okay. 1.8 Completed GHRS/2 2537 (Physical Conditions of Interstellar Gas) The High Resolution Spectrograph was used to make high signal to noise ratio data in the direction of standard star HD80007 to study the interfaces between the warm and cold regions of the interstellar medium. The observations were received in real time. There was a deflection calibration followed by a spiral search which found the target in the second group. The final flux count of the target was 19,827. 1.9 Initiated FOS/BL 3791 (Quasar Absorption Line Survey) The Faint Object Spectrograph was used to initiate absorption line survey images of the quasar PG1216+069. The observations are currently being executed, and will be analyzed later. 1.10 Initiated WFC 4713 (WF/PC Calibration: WFC Internal Delta Flats) The Wide Field Camera was used to initiate internal flat observations which will be used for the restoration of Earth flat images which were altered due to changes in quantum efficiency which occurred after WFPC decontamination operations. The observations are currently ongoing, and will be analyzed later. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 4 (Failure at 074/1300Z and 074/0702Z) Primary was acquired, secondary was not. Dim secondary guide star (variable). Scheduled Reacquisitions : 8 Successful : 8 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 11 Successful : 11 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOS/BL 3791 (Quasar Absorption Line Survey) 3.2 Complete WFC 4713 (WF/PC Calibration: WFC Internal Delta Flats) 3.3 Complete PC 4714 (WF/PC Calibration: PC Internal Delta Flats) 3.4 Complete WFC/PC 4109 (WF/PC Calibration: Rapid Internal Monitor) 3.5 Complete PC 4347 (WF/PC Calibration: Flood Retention Test) 3.6 Complete GHRS/2 1168 (Interstellar Carbon and Oxygen) 3.7 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.8 Complete PC 4140 (WF/PC Calibration: Rapid Internal Monitor, Part 2) 3.9 Complete GHRS/2 3496 (Ultraviolet Atlas of Sirius) 3.10 Complete FOC/96 4804 (UV Imaging of Nearby Galaxies wit FOC/96) 3.11 Initiate WFC 4658 (Determination of Extragalactic Distance Scale) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: FGS #2 Anomaly Investigation Further testing and analysis of the FGS #2 anomalous behavior are still ongoing. In the interim FGS #2 will not be used for observations. 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-